log7(x) + log7(x)-log7(3)=log7(12)
Answer by stanbon(75887)

log7(x) + log7(x)-log7(3)=log7(12)
----
log7(x^2) = log7(12)-log7(3)
---
log7(x^2) = log7(4)
x^2 = 4
x = 2
---------
Cheers,
Stan H.
==========